Circular No. 6392

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1915+105 E. Morgan and R. Remillard, Massachusetts Institute of Technology;

and J. Greiner, Max-Planck-Institut für Extraterrestrische Physik, Garching, communicate: “Recent RXTE observations have shown that GRS 1915+105, a galactic x-ray transient (IAUC 5590) with relativistic radio jets (Mirabel and Rodriguez 1994, Nature 371, 46), is bright and highly variable in the range 2–25 keV. More than 800 measurements from the All Sky Monitor (ASM) at 2–12 keV during 1996 Feb. 20–Apr. 26 show large intensity variation from 0.2 to 2.0 Crab. The ASM captured several episodes of quasiperiodic oscillations (QPOs) with timescales of 30–40 s and amplitudes as large as 1 Crab. Five PCA observations taken in April show many narrow QPO features between 0.5 and 100 mHz with complex spectral and dynamic signatures and large amplitudes. There is a trend toward smaller amplitude and higher frequency variability, although the average intensity has remained ∼ 1.5 Crab. In the two most recent PCA observations, we find a 67-Hz QPO. If this feature is related to the frequency of the minimum stable Keplerian orbit around a Schwartzschild blackhole, then the implied blackhole mass is 33 M_{⊙}. Radio Observations are urged to look for new episodes of superluminal mass ejection. RXTE will continue to monitor this source.”

SUPERNOVA 1996M IN ANONYMOUS GALAXY A. Filippenko, Univerisity of California, reports that SN 1996M is due

east of the host galaxy’s nucleus, rather than west (cf. IAUC 6347).

COMET 46P/WIRTANEN CCD R magnitude estimates obtained at the European Southern Ob-

servatory: 1995 June 26.23 UT, 24.2 (H. Böhnhardt, with R. M. West and J. Babion, 2.2-m reflector; stellar appearance); 1996 Apr. 25.33, 21.5 (Böhnhardt and H. Rauer, with S. Mottola and A. Nathues, 1.5-m Danish reflector; 4′′ tail in p.a. ∼ 270^{o}).

COMET 65P/GUNN Total visual magnitude estimates (cf. IAUC 6368): Apr. 15.44 UT,

13.3 (A. Hale, Cloudcroft, NM, 0.41-m reflector); 18.06, 13.0 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 20.09, 13.1 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 27.43, 12.9 (Hale).

1996 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green